# OpenXAI: Towards a Transparent Evaluation of Model Explanations

### v1

@@ 2 Overview of OpenXAI Framework
d) Fairness: Following the work by Dai et al. [12], we measure the fairness of post hoc explanations by averaging all the aforementioned metric values across instances in the majority and minority subgroups and comparing the two estimates. If there is a huge difference in the two estimates, then we consider this to be evidence for unfairness.
Detailed descriptions of all the aforementioned metrics along with notation and equations are included in the Appendix. Invoking the aforementioned metrics to benchmark an explanation methods is quite simple and the code snippet below describes how to invoke the RIS metric. Users can easily incorporate their own custom evaluation metrics into OpenXAI by filling a form and providing the GitHub link to their code as well as a summary of their metric (See Appendix). Note that the code should be in the form of a function which takes as input data instances, corresponding model predictions and their explanations, as well as OpenXAI’s model object and returns a numerical score. from openxai import Evaluatormetric_evaluator = Evaluator(inputs, labels, model, explanations)score = metric_evaluator.eval(metric=‘RIS’)
@@
Benchmarking: As can be seen from the code snippets in this section, OpenXAI allows end users to easily benchmark explanation methods using just a few lines of code. To summarize the benchmarking process, let us consider a scenario where we would like to benchmark a new explanation method using OpenXAI’s pre-trained neural network model and the German Credit dataset. First, we use OpenXAI’s Dataloader class to load the German Credit dataset. Second, we load the neural network model (’ann’) using our LoadModel class. Third, we extend the Explainer class and incorporate the code for the new explanation method in the get_explanation function of this class. Finally, we evaluate the new explanation method using various metrics from the Evaluator class.
4) Leaderboards. OpenXAI introduces the first ever public XAI leaderboards to promote transparency, and enable users to easily compare the performance of multiple explanation methods across a variety of evaluation metrics, predictive models, and datasets. In the current release, we have six different leaderboards each corresponding to a particular dataset. A snapshot of one of our leaderboard pages is shown in Figure 2. Users can submit requests for their custom explanation methods to be featured on one of our leaderboards. To this end, they first need to following the aforementioned benchmarking process to develop and evaluate their explanation method. Then, they will need to fill a simple form, and provide the GitHub link to their code, a summary of their explanation method, and report the average and standard error values for each of the metrics. We then verify the submitted code and corresponding results, and add the new method to our leaderboard upon successful verification.
Figure 2: A snapshot of the leaderboard page from OpenXAI public website. We also provide interactive ranking functionality (arrow mark in the figure) which allows users to rank explanation methods based on metrics of their choice. Please visit the website to see leaderboards for other datasets.
@@ 3 Benchmarking Analysis
Next, we describe how we benchmark state-of-the-art explanation methods using our OpenXAI framework, and also discuss key findings of this benchmarking analysis. Code to reproduce all the results is available at https://github.com/AI4LIFE-GROUP/OpenXAI.
Experimental Setup. We benchmark all the six state-of-the-art feature attribution methods currently available in our OpenXAI framework along with the random baseline, using the openxai.Evaluator module (See Section 2). We use default hyperparameter settings for all these methods following the guidelines outlined in the original implementations. Details about the hyperparameters used in our experiments are discussed in Section E.3 in the Appendix. Our OpenXAI framework currently has two pre-trained models, a logistic regression model and a deep neural network model, for each dataset. The neural network models have two fully connected hidden layers with 100 nodes in each layer, and they use ReLU activation functions and an output softmax layer. See Appendix E.4 for more details on model architectures, model training, and model performance.
Table 2: Ground-truth and predicted faithfulness results on the Heloc dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across all instances in the test set. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@
[…]
$2±\pm0.0$0
[…]
$2±\pm0.0$0
[…]
Table 3: Ground-truth and predicted faithfulness results on the Adult Income dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across all instances in the test set. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@
[…]
$\pm0.0$0
[…]
Ground-truth faithfulness: We evaluate ground-truth faithfulness by calculating the similarity between the generated explanations and the ground-truth explanations using the metrics discussed in Section 2. Results for various ground-truth faithfulness metrics are shown in Tables 2, 3, 8, 9. Vanilla Gradients, SmoothGrad, and Integrated Gradients produce explanations that achieve perfect scores on four ground-truth faithfulness metrics, viz. pairwise rank agreement (PRA), feature agreement (FA), rank agreement (RA), and rank correlation (RC) metrics, for all datasets. However, on average, across all datasets, LIME outperforms other methods on the signed agreement (SA) [+61.6%] and signed-rank agreement (SRA) [+65.3%] metrics, whereas gradient-based explainers achieve relatively lower values. While illustrative in nature, these findings show how OpenXAI can help identify the limitations of existing explanation methods, which in turn can inform the design of new methods.
Predictive faithfulness: Tables 2, 3, 8, 9 show results for the PGI and PGU metrics implemented in OpenXAI (see Section 2 and Appendix A). Overall, we find that SmoothGrad explanations are most faithful to the underlying model and, on average, across multiple datasets outperform other feature-attribution methods on PGU metric (+30.7%). However, the OpenXAI leaderboard for the HELOC dataset reveals that LIME produces considerably more faithful (+24.9%) explanations than other methods, identifying a setting where LIME approximates the underlying model more accurately. Finally, this analysis confirms the finding by Krishna et al. [34] that explanations output by state-of-the-art methods do not necessarily align with each other. This finding further highlights the need for rigorous empirical and theoretical benchmarking of explanation methods.
Stability. Next, we examine the stability of explanation methods when the underlying models are LR models in Tables 3 and 3, and neural network models in Tables E.5 and E.5 in the Appendix. Due to space constraints, we focus on RIS and RRS metrics in the main paper and leave the other results to the Appendix. Overall, the relative stability varies considerably across different datasets, implying that no single explanation method is consistently the most stable. First, for the synthetic dataset in Table 3, we find that Gradient x Input, on average, outperforms feature-attribution methods in relative input stability (+93.5%, RIS) and relative representation stability (+59.2%, RRS). However, stability of Gradient x Input significantly degrades on real-world datasets (Table 3, E.5, E.5). Second, as shown in Tables 3, E.5, and E.5, there is no single explanation method that has the highest input and representation stability across all the real-world datasets. On average, across all real-world datasets, SmoothGrad achieves 63.2% higher RRS values compared to other methods, whereas no method performs consistently well when it comes to the RIS metric.
Table 4: Stability of explanation methods on the Synthetic dataset with LR model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS
@@
Vanilla Gradients
Integrated Gradients
@@
[…]
Table 5: Stability of explanation methods on the German Credit dataset with LR model. Shown are average and standard error values across all test set instances. Values closer to zero are desirable, and the best performance is bolded.
Method RIS RRS
@@
Vanilla Gradients
Integrated Gradients
@@
[…]
Fairness. To measure fairness of explanation methods, we compute the average metric values (for each of the aforementioned faithfulness and stability metrics) for different subgroups (e.g., male and female) in the dataset and compare them. Without loss of generality, we present results using the PGU (see Section 2 and Appendix A) metric. Results for LR models in Figures 3 and 4 provide two key findings. First, the fairness analysis in Figures 3 and 4 shows that there are disparities in the faithfulness of explanations (see Section 2) output by several methods (Vanilla Gradients, Integrated Gradients, and SmoothGrad). Second, Gradient x Input results in the least amount of disparity across both the datasets. These results also suggest a trade-off between various evaluation metrics. For instance, Gradient x Input method underperforms on faithfulness and stability metrics, but outperforms other methods (+8.9%) when it comes to fairness metrics. Given such trade-offs, practitioners can leverage the OpenXAI leaderboards (Figure 2) to select an explanation method that best meets application-specific needs. Results with NN models and other fairness metrics are included in the Appendix E.6.
Figure 3: Fairness analysis of PGU metric on the German Credit dataset with LR model. Shown are average and standard error values for majority (male) and minority (female) subgroups.
Figure 4: Fairness analysis of PGU metric on the Adult Income dataset with LR model. Shown are average and standard error values for majority (male) and minority (female) subgroups.
4 Conclusions
@@ Appendix A Evaluation Metrics
1) Faithfulness. To measure how faithfully a given explanation mimics the underlying model, prior work has either leveraged synthetic datasets to obtain ground truth explanations or measured the differences in predictions when feature values are perturbed [48]. Here, we discuss the two broad categories of faithfulness metrics included in OpenXAI, namely ground-truth and predictive faithfulness.
@@
2) Stability. We leverage the metrics Relative Input Stability (RIS), Relative Representation Stability (RRS), and Relative Output Stability (ROS) [1] which measure the maximum change in explanation relative to changes in the inputs, internal representations learned by the model, and output prediction probabilities respectively. These metrics can be written formally as:
[…]
@@
[…]
@@
[…]
@@
3) Fairness. It is important to ensure that there are no significant disparities between the reliability of post hoc explanations corresponding to instances in the majority and the minority subgroups. To this end, we average all the aforementioned metric values across instances in the majority and minority subgroups, and then compare the two estimates (See Figures 4, 3, 6 etc.) to check if there are significant disparities [12].
@@ Appendix B Additional Related Work
Inherently Interpretable Models and Post hoc Explanations.
Many approaches learn inherently interpretable models such as rule lists [67, 64], decision trees and decision lists [39], and others [37, 8, 43, 9]. However, complex models such as deep neural networks often achieve higher accuracy than simpler models [51]. Thus, there has been significant interest in constructing post hoc explanations to understand their behavior. To this end, several techniques have been proposed in recent literature to construct post hoc explanations of complex decision models. For instance, LIME, SHAP, Anchors, BayesLIME, and BayesSHAP [51, 44, 52, 58] are considered perturbation-based local explanation methods because they leverage perturbations of individual instances to construct interpretable local approximations (e.g., linear models). On the other hand, methods such as Vanilla Gradients, Gradient x Input, SmoothGrad, Integrated Gradients, and GradCAM [56, 60, 54, 59] are referred to as gradient-based local explanation methods since they leverage gradients computed with respect to input features of individual instances to explain individual model predictions.
There has also been recent work on constructing counterfactual explanations which capture what changes need to be made to a given instance in order to flip its prediction [63, 61, 28, 50, 42, 6, 29, 46, 30, 47]. Such explanations can be leveraged to provide recourse to individuals negatively impacted by algorithmic decisions. An alternate class of methods referred to as global explanation methods attempt to summarize the behavior of black-box models as a whole rather than in relation to individual data points [38, 7]. A more detailed treatment of this topic is provided in other comprehensive survey articles [4, 20, 45, 40, 10].
@@
Evaluating Post hoc Explanations. In addition to the quantitative metrics designed to evaluate the reliability of post hoc explanation methods [41, 68, 3, 12] (See Section 1), prior works have also introduced human-grounded approaches to evaluate the interpretability of explanations generated by these methods [13]. For example, Lakkaraju and Bastani [36] carry out a user study to understand if misleading explanations can fool domain experts into deploying racially biased models, while Kaur et al. [31] find that explanations are often over-trusted and misused. Similarly, Poursabzi-Sangdeh et al. [49] find that supposedly-interpretable models can lead to a decreased ability to detect and correct model mistakes, possibly due to information overload. Jesus et al. [26] introduce a method to compare explanation methods based on how subject matter experts perform on specific tasks with the help of explanations. Lage et al. [35] use insights from rigorous human-subject experiments to inform regularizers used in explanation algorithms. In contrast to the aforementioned research, our work leverages twenty-two different state-of-the-art quantitative metrics to systematically benchmark the reliability (and not interpretability) of post hoc explanation methods.
Limitations and Vulnerabilities of Post hoc Explanations. Various quantitative metrics proposed in literature (See Section 1) were also leveraged to analyze the behavior of post hoc explanation methods and their vulnerabilities—e.g., Ghorbani et al. [19] and Slack et al. [57] demonstrated that methods such as LIME and SHAP may result in explanations that are not only inconsistent and unstable, but also prone to adversarial attacks. Furthermore Lakkaraju and Bastani [36] and Slack et al. [57] showed that explanations which do not accurately represent the importance of sensitive attributes (e.g., race, gender) could potentially mislead end users into believing that the underlying models are fair when they are not [36, 57, 2]. This, in turn, could lead to the deployment of unfair models in critical real world applications. There is also some discussion about whether models which are not inherently interpretable ought to be used in high-stakes decisions at all. Rudin [53] argues that post hoc explanations tend to be unfaithful to the model to the extent that their usefulness is severely compromised. While this line of work demonstrates different ways in which explanations could potentially induce inaccuracies and biases in real world applications, they do not focus on systematic benchmarking of post hoc explanation methods which is the main goal of our work.

### v5

@@ 2 Overview of OpenXAI Framework
d) Fairness: Following the work by Dai et al. (2022), we measure the fairness of post hoc explanations by averaging all the aforementioned metric values across instances in the majority and minority subgroups, and comparing the two estimates. If there is a huge difference in the two estimates, then we consider this to be evidence for unfairness.
Invoking the aforementioned metrics to benchmark an explanation methods is quite simple and the code snippet below describes how to invoke the PGI metric (where kwargs represents a dictionary containing the given metric’s parameters).
from openxai.evaluator import Evaluatormetric_evaluator = Evaluator(model, metric=‘PGI’)score, mean_score = metric_evaluator.evaluate(**kwargs)
@@
Benchmarking: As can be seen from the code snippets in this section, OpenXAI allows end users to easily benchmark explanation methods using just a few lines of code. To summarize the benchmarking process, let us consider a scenario where we would like to benchmark a new explanation method using OpenXAI’s pre-trained neural network model and the German Credit dataset. First, we use OpenXAI’s ReturnLoaders function to load the German Credit dataset. Second, we load the neural network model (‘ann’) using our LoadModel function. Third, we extend the Explainer class and incorporate the code for the new explanation method in the get_explanations function of this class. Finally, we evaluate the new explanation method using any metric from the Evaluator class.
@@ 3 Benchmarking Analysis
Next, we describe how we benchmark state-of-the-art explanation methods using our OpenXAI framework and also discuss the key findings of this benchmarking analysis. Code to reproduce all the results is available at https://github.com/AI4LIFE-GROUP/OpenXAI.
Experimental Setup. We benchmark all of the six state-of-the-art feature attribution methods currently available in our OpenXAI framework along with the random baseline, using the openxai.Evaluator module (See Section 2), and for the first 1000 test instances of each dataset. Details about the hyperparameters used in our experiments are discussed in Section  D.3 in the Appendix. Our OpenXAI framework currently has two pre-trained models, a logistic regression model and a deep neural network model, for each dataset. The neural network models have two fully connected hidden layers with 100 nodes in each layer, and they use ReLU activation functions and an output softmax layer. See Appendix D.4 for details on model architectures, training and performance.
Ground-truth faithfulness: For logistic regression models, we evaluate ground-truth faithfulness by calculating the similarity between the generated explanations and the ground-truth explanations using the metrics discussed in Section 2. Results for various ground-truth faithfulness metrics are shown in Tables 2-3, 8-9, 16-19. Vanilla Gradients, Integrated Gradients, and SmoothGrad produce explanations that achieve perfect scores on four ground-truth faithfulness metrics, viz. pairwise rank agreement (PRA), rank correlation (RC), feature agreement (FA), rank agreement (RA), sign agreement (SA), signed rank agreement (SRA) metrics, for all datasets. LIME outperforms other methods, approaching the performance of the above gradient-based explanations in many cases. While illustrative in nature, these findings show how OpenXAI can help identify the limitations of existing explanation methods, which in turn can inform the design of new methods.
Predictive faithfulness: Tables 2-3, 8-9, 16-19 show results for logistic regression models, and Tables D.5-D.5, D.7-D.7 show results for neural network models on the PGI and PGU metrics implemented in OpenXAI (see Section 2 and Appendix A). Overall, we find that Vanilla Gradients, SmoothGrad, and LIME explanations are most faithful to the underlying model and, on average across all datasets and models, outperform other feature-attribution methods on the PGI metric. The PGU metric showed similar trends in terms of best performance, with Integrated Gradients on par with the leading metrics.
Table 2: Ground-truth and predictive faithfulness results on the HELOC dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@
0.500$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.599$\pm$0.00
1.000$\pm$0.00
0.603$\pm$0.00
0.923$\pm$0.00
0.002$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.291$\pm$0.01
1.000$\pm$0.00
0.291$\pm$0.01
0.946$\pm$0.00
0.149$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.205$\pm$0.00
1.000$\pm$0.00
0.195$\pm$0.00
0.910$\pm$0.00
0.046$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.029$\pm$0.00
1.000$\pm$0.00
0.029$\pm$0.00
0.682$\pm$0.01
0.076$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.205$\pm$0.00
1.000$\pm$0.00
0.195$\pm$0.00
0.910$\pm$0.00
0.024$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.029$\pm$0.00
1.000$\pm$0.00
0.029$\pm$0.00
0.682$\pm$0.01
0.098$\pm$0.00
0.073$\pm$0.00
0.073$\pm$0.00
0.096$\pm$0.00
0.073$\pm$0.00
0.096$\pm$0.00
0.073$\pm$0.00
0.039$\pm$0.00
0.076$\pm$0.00
0.076$\pm$0.00
0.046$\pm$0.00
0.076$\pm$0.00
0.045$\pm$0.00
0.076$\pm$0.00
Table 3: Ground-truth and predictive faithfulness results on the Adult Income dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@
0.501$\pm$0.00
1.000$\pm$0.00
1.000$\pm$0.00
0.538$\pm$0.00
1.000$\pm$0.00
0.627$\pm$0.00
0.881$\pm$0.00
0.003$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.215$\pm$0.00
1.000$\pm$0.00
0.319$\pm$0.01
0.876$\pm$0.00
0.188$\pm$0.01
0.999$\pm$0.00
1.000$\pm$0.00
0.359$\pm$0.00
0.999$\pm$0.00
0.363$\pm$0.00
0.924$\pm$0.00
0.073$\pm$0.00
0.998$\pm$0.00
1.000$\pm$0.00
0.138$\pm$0.01
0.999$\pm$0.00
0.041$\pm$0.00
0.875$\pm$0.00
0.093$\pm$0.00
0.999$\pm$0.00
1.000$\pm$0.00
0.359$\pm$0.00
0.999$\pm$0.00
0.362$\pm$0.00
0.924$\pm$0.00
0.034$\pm$0.00
0.998$\pm$0.00
1.000$\pm$0.00
0.138$\pm$0.01
0.999$\pm$0.00
0.041$\pm$0.00
0.875$\pm$0.00
0.130$\pm$0.00
0.065$\pm$0.00
0.065$\pm$0.00
0.124$\pm$0.00
0.065$\pm$0.00
0.125$\pm$0.00
0.063$\pm$0.00
0.043$\pm$0.00
0.124$\pm$0.00
0.124$\pm$0.00
0.062$\pm$0.00
0.124$\pm$0.00
0.062$\pm$0.00
0.126$\pm$0.00
Stability. Next, we examine the stability of explanation methods when the underlying models are logistic regression models in Tables  3-3 below (additionally, Tables D.5-D.5, D.6D.6 in the Appendix), and neural network models (Tables D.5-D.5, D.7D.7). For logistic regression models, we consider only the RIS and ROS metrics for stability, since RRS is equivalent to ROS within this model class. Overall, we see similar trends across explanation methods for each of the relative stability metrics at our chosen set of hyperparameters (see Appendix D.3). First, Integrated Gradients tends to outperform all other feature-attribution methods with log average scores of 1.95 and 1.86 for RIS and ROS, respectively, when averaged across all datasets. Vanilla Gradients and Gradient x Input remain on roughly the same order of magnitude with log average scores of 3.16 and 3.37 for RIS, and 2.91 and 4.49 for ROS. The remaining methods, LIME, SHAP, SmoothGrad and Random, score log RIS and ROS scores between roughly 10 and 15.
Table 4: Stability results on the Synthetic dataset for all explanation methods with LR model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) ROS ($\downarrow$)
@@
VanillaGrad
IntegratedGrad
@@
11.58$\pm$0.00
2.49$\pm$0.01
1.17$\pm$0.01
2.50$\pm$0.01
9.69$\pm$0.01
9.35$\pm$0.01
9.10$\pm$0.02
12.34$\pm$0.02
0.85$\pm$0.04
-0.70$\pm$0.02
2.82$\pm$0.04
11.25$\pm$0.02
11.69$\pm$0.02
12.17$\pm$0.04
Table 5: Stability results on the German Credit dataset for all explanation methods with LR model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) ROS ($\downarrow$)
@@
VanillaGrad
IntegratedGrad
@@
12.84$\pm$0.00
-0.01$\pm$0.04
-0.78$\pm$0.03
0.47$\pm$0.01
8.87$\pm$0.03
10.01$\pm$0.07
9.59$\pm$0.01
14.78$\pm$0.04
-0.21$\pm$0.07
-1.20$\pm$0.05
3.10$\pm$0.09
10.71$\pm$0.07
13.88$\pm$0.10
11.34$\pm$0.05
Fairness. To measure fairness of explanation methods, we compute the average metric values (for each of the aforementioned faithfulness and stability metrics) for different subgroups (e.g., male and female) in the dataset and compare them. Larger gaps between the metric values for different subgroups indicates higher disparities (unfairness). We present results for the PGU (see Section 2 and Appendix A) metric. The fairness analysis in Figures  1 and 2 shows that there are disparities in the faithfulness of explanations (see Section 2) output by several methods (most notably, Gradient x Input and SHAP on the Adult Income dataset). Results with NN models are included in Appendix D.7.
[Uncaptioned image] Figure 1: Fairness analysis of PGU metric on the German Credit dataset with LR model. Shown are average and standard error values for majority (male) and minority (female) subgroups. Larger gaps between the values of majority and minority subgroups (i.e., red and blue bars respectively) indicate higher disparities which are undesirable. [Uncaptioned image] Figure 2: Fairness analysis of PGU metric on the Adult Income dataset with LR model. Shown are average and standard error values for the majority (male) and minority (female) subgroups. Larger gaps between the values of majority and minority subgroups (i.e., red and blue bars respectively) indicate higher disparities which are undesirable.
4 Conclusion
Acknowledgments and Disclosure of Funding
@@ Appendix A Evaluation Metrics
1) Faithfulness. To measure how faithfully a given explanation mimics the underlying model, prior work has either leveraged synthetic datasets to obtain ground truth explanations or measured the differences in predictions when feature values are perturbed [Petsiuk et al., 2018]. Here, we discuss the two broad categories of faithfulness metrics included in OpenXAI, namely ground-truth and predictive faithfulness.
@@
2) Stability. We leverage the metrics Relative Input Stability (RIS), Relative Representation Stability (RRS), and Relative Output Stability (ROS) Agarwal et al. [2022] which measure the maximum change in explanation relative to changes in the inputs, internal representations learned by the model, and output prediction probabilities respectively. These metrics can be written formally as:
$\textup{RIS}=\frac{||\mathbf{x}||_{p}}{||\mathbf{e}_{\mathbf{x}}||_{p}}\max_{\mathbf{x}^{\prime}}\frac{||\mathbf{e}_{\mathbf{x}}-\mathbf{e}_{\mathbf{x}^{\prime}}||_{p}}{||\mathbf{x}-\mathbf{x}^{\prime}||_{p}},\leavevmode\nobreak\ \forall\mathbf{x}^{\prime}\leavevmode\nobreak\ \textup{s.t.}\leavevmode\nobreak\ \mathbf{x}^{\prime}\in\mathcal{N}_{\mathbf{x}};\leavevmode\nobreak\ \hat{y}_{\mathbf{x}}=\hat{y}_{\mathbf{x}^{\prime}}$ (3)
@@
$\textup{RRS}=\frac{||\mathcal{L}_{\mathbf{x}}||_{p}}{||\mathbf{e}_{\mathbf{x}}||_{p}}\max_{\mathbf{x}^{\prime}}\frac{{||\mathbf{e}_{\mathbf{x}}-\mathbf{e}_{\mathbf{x}^{\prime}}||_{p}}}{||\mathcal{L}_{\mathbf{x}}{-}\mathcal{L}_{\mathbf{x}^{\prime}}||_{p}},\leavevmode\nobreak\ \forall\mathbf{x}^{\prime}\leavevmode\nobreak\ \leavevmode\nobreak\ \textup{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{x}^{\prime}\in\mathcal{N}_{\mathbf{x}};\leavevmode\nobreak\ \leavevmode\nobreak\ \hat{y}_{\mathbf{x}}{=}\hat{y}_{\mathbf{x}^{\prime}}$ (4)
@@
$\textup{ROS}=\frac{||f({\mathbf{x}})||_{p}}{||\mathbf{e}_{\mathbf{x}}||_{p}}\max_{\mathbf{x}^{\prime}}\frac{||\mathbf{e}_{\mathbf{x}}-\mathbf{e}_{\mathbf{x}^{\prime}}||_{p}}{||f({\mathbf{x}}){-}f({\mathbf{x}^{\prime}})||_{p}},\leavevmode\nobreak\ \forall\mathbf{x}^{\prime}\leavevmode\nobreak\ \leavevmode\nobreak\ \textup{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{x}^{\prime}\in\mathcal{N}_{\mathbf{x}};\leavevmode\nobreak\ \leavevmode\nobreak\ \hat{y}_{\mathbf{x}}{=}\hat{y}_{\mathbf{x}^{\prime}},$ (5)
@@
3) Fairness. It is important to ensure that there are no significant disparities between the reliability of post hoc explanations corresponding to instances in the majority and the minority subgroups. To this end, we average all the aforementioned metric values across instances in the majority and minority subgroups, and then compare the two estimates to check if there are significant disparities Dai et al. [2022]. See Figures 1 to 2 for logistic regression models, and Figures 3 to 4 for neural network models.
@@ Appendix B Additional Related Work
Inherently Interpretable Models and Post hoc Explanations. Many approaches learn inherently interpretable models such as rule lists [Zeng et al., 2017, Wang and Rudin, 2015], decision trees and decision lists [Letham et al., 2015], and others [Lakkaraju et al., 2016, Bien and Tibshirani, 2009, Lou et al., 2012, Caruana et al., 2015]. However, complex models such as deep neural networks often achieve higher accuracy than simpler models Ribeiro et al. [2016]. Thus, there has been significant interest in constructing post hoc explanations to understand their behavior. To this end, several techniques have been proposed in recent literature to construct post hoc explanations of complex decision models. For instance, LIME, SHAP, Anchors, BayesLIME, and BayesSHAP Ribeiro et al. [2016], Lundberg and Lee [2017b], Ribeiro et al. [2018], Slack et al. [2021] are considered perturbation-based local explanation methods because they leverage perturbations of individual instances to construct interpretable local approximations (e.g., linear models). On the other hand, methods such as Vanilla Gradients, Gradient x Input, SmoothGrad, Integrated Gradients, and GradCAM Simonyan et al. [2014], Sundararajan et al. [2017], Selvaraju et al. [2017], Smilkov et al. [2017] are referred to as gradient-based local explanation methods since they leverage gradients computed with respect to input features of individual instances to explain individual model predictions.
There has also been recent work on constructing counterfactual explanations which capture what changes need to be made to a given instance in order to flip its prediction Wachter et al. [2017], Ustun et al. [2019], Karimi et al. [2019], Poyiadzi et al. [2020], Looveren and Klaise [2019], Barocas et al. [2020], Karimi et al. [2020a], Pawelczyk et al. [2020], Karimi et al. [2020b], Pawelczyk et al. [2021]. Such explanations can be leveraged to provide recourse to individuals negatively impacted by algorithmic decisions. An alternate class of methods referred to as global explanation methods attempt to summarize the behavior of black-box models as a whole rather than in relation to individual data points [Lakkaraju et al., 2019, Bastani et al., 2017]. A more detailed treatment of this topic is provided in other comprehensive survey articles [Arrieta et al., 2020, Guidotti et al., 2018, Murdoch et al., 2019, Linardatos et al., 2021, Covert et al., 2021].
@@
Evaluating Post hoc Explanations. In addition to the quantitative metrics designed to evaluate the reliability of post hoc explanation methods Liu et al. [2021], Zhou et al. [2021], Alvarez-Melis and Jaakkola [2018], Dai et al. [2022] (See Section 1), prior works have also introduced human-grounded approaches to evaluate the interpretability of explanations generated by these methods [Doshi-Velez and Kim, 2017]. For example, Lakkaraju and Bastani [2020] carry out a user study to understand if misleading explanations can fool domain experts into deploying racially biased models, while Kaur et al. [2020] find that explanations are often over-trusted and misused. Similarly, Poursabzi-Sangdeh et al. [2018] find that supposedly-interpretable models can lead to a decreased ability to detect and correct model mistakes, possibly due to information overload. Jesus et al. [2021] introduce a method to compare explanation methods based on how subject matter experts perform on specific tasks with the help of explanations. Lage et al. [2019] use insights from rigorous human-subject experiments to inform regularizers used in explanation algorithms. In contrast to the aforementioned research, our work leverages twenty-two different state-of-the-art quantitative metrics to systematically benchmark the reliability (and not interpretability) of post hoc explanation methods.
Limitations and Vulnerabilities of Post hoc Explanations. Various quantitative metrics proposed in literature (See Section 1) were also leveraged to analyze the behavior of post hoc explanation methods and their vulnerabilities—e.g., Ghorbani et al. [2019] and Slack et al. [2020] demonstrated that methods such as LIME and SHAP may result in explanations that are not only inconsistent and unstable, but also prone to adversarial attacks. Furthermore Lakkaraju and Bastani [2020] and Slack et al. [2020] showed that explanations which do not accurately represent the importance of sensitive attributes (e.g., race, gender) could potentially mislead end users into believing that the underlying models are fair when they are not Lakkaraju and Bastani [2020], Slack et al. [2020], Aivodji et al. [2019]. This, in turn, could lead to the deployment of unfair models in critical real world applications. There is also some discussion about whether models which are not inherently interpretable ought to be used in high-stakes decisions at all. Rudin [2019] argues that post hoc explanations tend to be unfaithful to the model to the extent that their usefulness is severely compromised. While this line of work demonstrates different ways in which explanations could potentially induce inaccuracies and biases in real world applications, they do not focus on systematic benchmarking of post hoc explanation methods which is the main goal of our work.